You’re standing in waist-deep water with your eyes 0.46 m m above the surface. A fish is swimming 0.65 m below the surface. Your line of sight to the fish makes a 38 ∘ angle with the water’s surface.

How far (horizontally) from you is the fish?

Express your answer to two significant figures and include the appropriate units.
Answer #1

Angle of incidence i = 90 - 38 = 52 degree

angle of reflection r = arcsin ((sin i)/n) = arcsin ((sin 52 degree)/1.33) = 36.33 degree

Horizontal Distance = 0.46*tan 52 degree + 0.65 tan 36.33 degree

= 1.1 m answer